Imagine you have 8 colored balls in a bag: 2 yellow, 2 red, 2 blue, and 2 green. If you pull out 4 balls, what is the probability that you have pulled out one of each color?

Answer by sudhanshu_kmr(1152)


no. of ways to pull 4 balls from 8 = 8C4
no. of ways to pull one ball from each color = 2C1 * 2C1 * 2C1 *2C1 = 16

probability = 16/8C4 = 16 /70 = 8/35